OpenStudy (anonymous):

You are planning to use a ceramic tile design in your new bathroom. The tiles are blue-and-white equilateral triangles. You decide to arrange the blue tiles in a hexagonal shape as shown. if the side of each tile measures 3 centimeters, what will be the exact area of each hexagonal shape? a. 9 cm b. 18 (square root) 3 cm c. 13.5 (square root) 3 cm d. 81 cm

OpenStudy (raden):

a hexagonal equals 6 times area of a triangle. and to get area of triangle you can use the formula 1/2 * s^2 * sqrt(3) (s = side of the equilateral triangle)

OpenStudy (raden):

so, 6 * 1/2 * s^2 * sqrt(3) put s = 3, get 6 * 1/2 * 3^2 * sqrt(3) = ...

OpenStudy (anonymous):

what does the ^ mean?

OpenStudy (raden):

^ = power = exponent it like 2^3 = 2 * 2 * 2 = 8

OpenStudy (anonymous):

Thats what I thought

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (raden):

did you get it ? you;re welcome

OpenStudy (anonymous):

Kinda.